### v1

@@ 6.3 Model Evaluation: ACC Vehicle with Maximum Setting
Table 3: Root mean squared errors for different models used as car-following controller with maximum gap setting.
@@
[…]
Speed (m/s) 0.081 1.839 372817467.021 0.268
Spacing (m) 0.394 22.761 993362922.528 2.573
The testing models’ performance dynamics under the maximum gap ACC setting are different from those observed for the minimum gap setting. The RACER model maintains consistent efficacy, replicating its minimum gap setting proficiency as indicated in Table 3. The RACER model surpasses other models, especially in spacing and speed metrics. In particular, the performance of the NN model drastically declines in the maximum gap setting scenario, resulting in collisions during the simulation. While the NN model’s predictions using testing data align with training data, as depicted in Figure 8, its application as a CF controller results in erratic behavior, demonstrated by Figure 8. Such discrepancies highlight that a model’s performance during standard testing does not necessarily predict its real-world operational success, especially in this study of car-following modeling, it’s imperative to test the model by simulating driving behavior with the trained model, rather than solely relying on acceleration predictions from input data. Given the NN model’s collision incidents and subsequent simulation failure, it will be excluded from the following discussion. The RMSE for the NN model diverges significantly from the other models, indicating a disparity in predictive fidelity.
For the maximum gap setting, the OVRV model’s acceleration predictions have the highest RMSE (0.614 $m/s^{2}$), significantly exceeding the lowest recorded RMSE of 0.221 $m/s^{2}$ by the RACER model. The RMSE of the PINN model stands at a competitive 0.227 $m/s^{2}$, closely trailing RACER and demonstrating consistent performance across various driving conditions, including acceleration and deceleration phases, as shown in Figure 9. Nonetheless, the speed and spacing predictions by both the PINN and OVRV models do not match the accuracy of the RACER model, with the spacing RMSE being approximately an order of magnitude greater and the speed RMSE at least twice as large. This discrepancy is a testament to the challenges posed by maximum gap setting conditions and suggests that model efficacy during training does not always translate to dependable control capabilities.
@@
As presented from Figure 9 through 11, the OVRV model exhibits comparatively poor performance, particularly at the onset of the simulation. Notably, the spacing maintained by the OVRV model is less than that of both the RACER model and the originally observed spacing, raising potential safety concerns if used as a vehicle controller. The PINN model performs better than the OVRV model but still shows a divergence in the later stages of the simulation. This divergence contributes to its elevated RMSE in both speed and spacing metrics.
@@ 6.4 Model Evaluation: ACC Vehicle with Smoother Acceleration
In prior experiments, the actual vehicle acceleration was calculated using 0.1-second intervals, which introduced significant noise, as indicated by the high variance in the measurements. This noise obscured the assessment of our proposed RACER model’s response to acceleration oscillations, challenging the verification of the model’s rationality. To address this, we adopted a larger time interval of 0.5 seconds for a smoother estimation of the vehicle’s average acceleration over that that longer time step with the minimum gap setting. Figure 15 below illustrates the performance of various models. Notably, between 20 and 25 seconds, the RACER model uniquely adapts to the fluctuations in acceleration, influenced by the RDCs active during this period. Overall, the RACER model consistently surpasses competing models across all evaluated metrics, including acceleration, spacing, and speed as presented in Table 4.
@@
In addition to the performance metrics discussed, the training process of the RACER model incorporates the minimization of the RDCs loss, which is crucial for maintaining safe and stable car-following behavior. Figure 16 presents the RDCs loss progression throughout the training epochs. The loss is segmented into three components: speed loss, spacing loss, and relative speed loss. These components are critical as they correspond to the key aspects of rational car-following dynamics that our model aims to optimize.
@@ 7 Discussion & Conclusion
Our experimental findings highlight the significant differences between pure neural network models and their physically constrained counterparts, particularly with respect to handling nonlinear problems and offering predictions that surpass those provided by conventional models like OVRV. These results emphasize the critical importance of integrating physical constraints into neural networks used for acceleration predictions to ensure realistic long-term forecasts.
@@
Incorporating physical constraints into neural network models thus emerges as a valuable strategy, especially in accurately simulating the physical world. This approach holds significant promise for enhancing safety measures in the field of transportation.
We have not yet evaluated our model in the context of human driving data. Human driving behavior is undoubtedly more complex and diverse occasionally not strictly adhering to the RDCs. However, our proposed model, which effectively combines physical-guided AI with car-following principles, exhibits enormous potential to inform and guide human drivers, thus making driving safer and more rational. The model also holds promise for testing across a wide range of ACC vehicle driving scenarios as well as human driving situations. Future research could explore this exciting avenue and contribute to the progressive journey toward achieving safer, smarter, and more efficient transportation systems.

### v2

@@ 6.3 Model Evaluation: ACC Vehicle with Maximum Setting
Table 4: Root mean squared errors for different models used as car-following controllers with maximum gap setting.
@@
Acceleration (m/s2) 0.221 0.614 N/A (crash) 0.227
Speed (m/s) 0.081 1.839 N/A (crash) 0.268
Spacing (m) 0.394 22.761 N/A (crash) 2.573
The performance of the tested models under the maximum gap ACC setting diverges notably from that observed under the minimum gap setting. The RACER model maintains consistent efficacy, replicating its minimum gap setting proficiency as indicated in Table 4. The RACER model surpasses other models, especially in spacing and speed metrics. In particular, the performance of the NN model drastically declines in the maximum gap setting scenario, resulting in collisions during the simulation. In contrast, the NN model experiences a significant decline in performance under the maximum gap scenario, resulting in collisions during the simulation. While the NN model’s predictions align well with the training data, as shown in Fig. 9, its application as a car-following controller leads to erratic and unstable behavior, as depicted in Fig. 9. This highlights a key insight: a model’s strong performance in standard testing does not necessarily translate to success in real-world scenarios. Specifically, in the context of car-following modeling, it is critical to validate a model through driving behavior simulations rather than relying solely on acceleration predictions based on test data. Given the NN model’s collision incidents and subsequent simulation failure, it will be excluded from the following discussion. The RMSE for the NN model diverges significantly from the other models, indicating a disparity in predictive fidelity.
For the maximum gap setting, the OVRV model’s acceleration predictions have the highest RMSE (0.614 $m/s^{2}$), significantly exceeding the lowest recorded RMSE of 0.221 $m/s^{2}$ by the RACER model. The RMSE of the PINN model stands at a competitive 0.227 $m/s^{2}$, closely trailing RACER and demonstrating consistent performance across various driving conditions, including acceleration and deceleration phases, as shown in Fig. 10. Nonetheless, the speed and spacing predictions by both the PINN and OVRV models do not match the accuracy of the RACER model, with the spacing RMSE being approximately an order of magnitude greater and the speed RMSE at least twice as large. This discrepancy is a testament to the challenges posed by maximum gap setting conditions and suggests that model efficacy during training does not always translate to dependable control capabilities.
@@
As presented from Fig. 10 through 12, the OVRV model exhibits comparatively poor performance, particularly at the onset of the simulation. Notably, the spacing maintained by the OVRV model is less than that of both the RACER model and the originally observed spacing, raising potential safety concerns if used as a vehicle controller. The PINN model performs better than the OVRV model but still shows a divergence in the later stages of the simulation. This divergence contributes to its elevated RMSE in both speed and spacing metrics.
Fig. 13 to Fig. 15 showcase the RDC violations for the ACC vehicle with the maximum setting. Consistent with the earlier results from the minimum setting comparison, the proposed RACER model is the only model that fully complies with the RDCs. In contrast, both the baseline NN and PINN models exhibit violations in one or more of the constraints. Specifically, the NN model fails in this scenario, leading to crashes and irrational predictions. These results further emphasize the advantage of incorporating RDCs into the learning process, as demonstrated by the superior performance of the RACER model.
@@ 6.4 Model Evaluation: ACC Vehicle with Smoother Acceleration
In prior experiments, the actual vehicle acceleration was calculated using 0.1-second intervals, which introduced significant noise, as indicated by the high variance in the measurements. This noise obscured the assessment of our proposed RACER model’s response to acceleration oscillations, challenging the verification of the model’s rationality. To address this, we adopted a larger time interval of 0.5 seconds for a smoother estimation of the vehicle’s average acceleration over that that longer time step with the minimum gap setting. Fig. 16 below illustrates the performance of various models. Notably, between 20 and 25 seconds, the RACER model uniquely adapts to the fluctuations in acceleration, influenced by the RDCs active during this period. Overall, the RACER model consistently surpasses competing models across all evaluated metrics, including acceleration, spacing, and speed as presented in Table 5.
@@
In addition to the performance metrics discussed, the training process of the RACER model incorporates the minimization of the RDCs loss, which is crucial for maintaining safe and stable car-following behavior. Fig. 17 presents the RDCs loss progression throughout the training epochs. The loss is segmented into three components: speed loss, spacing loss, and relative speed loss. These components are critical as they correspond to the key aspects of rational car-following dynamics that our model aims to optimize.
@@
As Table 6 shows, the RACER model uniquely combines the predictive power of deep learning with the hard safety bounds imposed by Rational Driving Constraints. Unlike the OVRV model, RACER captures complex driving patterns; unlike the NN model, it ensures physically meaningful outputs; and unlike the PINN model, it achieves strict RDC compliance rather than relying on a soft physics residual. This combination of high accuracy and provable driving rationality is particularly valuable for ACC vehicles and other safety-critical applications, where even occasional irrational predictions could have hazardous consequences. The present study therefore represents an initial step toward seamlessly incorporating domain-specific knowledge into AI-based car-following models.
Table 6: Comparison of Car-Following Modeling Approaches.
Characteristic Physics-Based Models Neural Networks PINN RACER
Foundation Mathematical equations with physical principles Data-driven learning from trajectories Hybrid approach combining NN with physics models NN with rational driving constraints
Physics Integration Hard-coded in model structure None (learned implicitly) Physics loss term in training Derivative constraints in loss
RDC Compliance Directly calculable through PDEs of explicit equations Not guaranteed Not directly enforced Explicitly enforced
RDC Performance Theoretically verifiable but limited model flexibility High accuracy but potential violations Good accuracy with some violations High accuracy with zero violations
Safety Guarantees Theoretical through model structure None Partial through physics guidance Complete through RDC enforcement
@@ 7 Discussion & Conclusion
Our experimental findings highlight the significant differences between pure neural network models and their physically constrained counterparts, particularly with respect to handling nonlinear problems and offering predictions that surpass those provided by conventional models like OVRV. These results emphasize the critical importance of integrating physical constraints into neural networks used for acceleration predictions to ensure realistic long-term forecasts. In contrast to existing approaches, our hybrid model bridges the gap between engineering and data-driven methods, combining the reliability of physical constraints with the adaptability of machine learning. This hybrid approach enables the model to maintain the integrity of real-world driving principles while leveraging data to optimize its predictions. While recent literature, including studies like [52, 53]), has emphasized the calibration of car-following models based on spacing, our approach introduces a novel perspective by focusing on acceleration as the control output. This decision is grounded in real-world applications, where controllers rely on acceleration or velocity rather than spacing to dictate vehicle behavior. As such, our model’s calibration aligns with the control dynamics observed in real-world ACC-equipped vehicles, ensuring a more practical and reliable implementation.
While the PINN framework has proven effective in implicitly capturing some of the outcomes of RDCs, our approach offers several key advancements by explicitly incorporating RDCs into the model’s loss function. This explicit integration ensures that the model adheres to these constraints throughout the learning process, resulting in more reliable and interpretable predictions. The explicit incorporation of RDCs, as demonstrated in our RACER model, offers enhanced flexibility by allowing the model to learn complex patterns from real-world driving data while maintaining adherence to fundamental driving constraints. This capability is essential for accurately capturing the nuanced behaviors observed in real-world driving scenarios—behaviors that purely physics-based models may struggle to represent due to their simplified mathematical formulations. The zero RDC violations achieved by RACER result from the hard constraint enforcement mechanism where gradient-based penalties immediately correct any tendency toward irrational behavior during training. In contrast, other models exhibit systematic violation patterns: NN models violate spacing constraints when trying to minimize prediction error in dense traffic scenarios, while PINN models trade off between data fitting and physics compliance, leading to violations when these objectives conflict.
@@
Incorporating physical constraints into neural network models has value, especially in accurately simulating the physical world. Our proposed controller aims to ensure the rationality of driving behavior. However, we do not assume that human driving behavior is always rational. This is why we focused our testing on ACC-equipped vehicles, which are automated vehicles that we assume should adhere to rational driving principles. The calibration of the RACER model relies on data from controlled experiments (Gunter et al. [11]), where the leading vehicle’s trajectory was pre-planned to elicit specific car-following patterns. While this approach enabled precise evaluation of the model under controlled conditions, it does not fully reflect the complexity of naturalistic driving scenarios. In real-world settings, ACC systems must respond to a variety of unpredictable behaviors, such as sudden braking when approaching a queue, lane changes by surrounding vehicles, or gaps created by diverging vehicles. These scenario-dependent factors may affect the ACC’s performance in ways not accounted for in the current dataset. Consequently, while the model performs well within the scope of the tested conditions, its generalizability to more diverse real-world scenarios remains to be fully assessed. Therefore, the RMSE values reported in the Results Section should be regarded as best-case performance under controlled conditions.
Human driving behavior is undoubtedly more complex and drivers occasionally (and often briefly) deviate from the RDCs. However, our proposed model, which effectively combines physical-guided AI with car-following principles, exhibits potential to inform and guide human drivers, thus making driving safer and more rational. The model also holds promise for testing across a wide range of ACC vehicle driving scenarios as well as human driving situations. To enhance the robustness and applicability of the RACER model, future research should prioritize the collection and integration of naturalistic ACC car-following data. Such data would expose the model to a wider spectrum of driving behaviors, enabling a more comprehensive evaluation and refinement of its performance across varied real-world conditions. This includes scenarios involving unpredictable lead vehicle behavior, lane changes, and other complex traffic interactions that are essential for full real-world deployment. Future research could explore this exciting avenue and contribute to the progressive journey toward achieving safer, smarter, and more efficient transportation systems.